a. [OH−]=0.002M b. [OH−]=1.2×10−4M c. [H3O+]=8.5×10−2M d. [H3O+]=4.2×10−8M Determine the pH for each

a. [OH−]=0.002M

b. [OH−]=1.2×10−4M

c. [H3O+]=8.5×10−2M

d. [H3O+]=4.2×10−8M

Determine the pH for each
Answer #1

pH = -log [H₂ot] pot = -log (oh) At 25c kw = 16-14 (ionic product of water ) Kw = [H₂ot] x [04] = 10-19 at 25° If pH <7 then(OH) = 0.002M poH = -log(0H] = -log(0.002) = 2.69897. pH = 14- pH = 14-2.69897= 11:3010 (OH)= 1.2 x 10 po -19 (1:2 X 10^4) =